A = Q, (a*b) = a + b


1
Expert's answer
2021-05-10T13:03:34-0400

"A=Q\left(a\cdot b\right)=a+b"


In the above expression, the identity element of the binary operation is zero.

From the identity's explanation e is the identity element if; "a\cdot e=a=e\cdot a"


The binary operation in our case will be: "a\cdot b=a+b-ab"

Letting e to be the binary element we get :

"a\cdot e=a\:\:"

"a+e-ae=a"

By subtracting a from both sides we get:

"e-ae=0\:"

"e\left(1-a\right)=0"


"a=1\:and\:e=0"


Therefore, the value of a can not be equal 1 because the binary operation * defined on Q-( 1 ).